OpenStudy (anonymous):

Find the diagonal of a cube if its side equals 5.

OpenStudy (dumbcow):

\[=\sqrt{5^{2}+ (5\sqrt{2})^{2}}=\sqrt{125}=5\sqrt{5}\]

OpenStudy (anonymous):

i think you're wrong dumbcow :) ?( if a side is 5 on the square, then you can use a² + b² = c² which would be 5² + 5² = 50 or sqrt(50) is your answer... because squares' internal angles are made of 90 degrees each... so you can use that Pythagorean formula

OpenStudy (dumbcow):

@paul you are correct for finding diagonal for square sqrt(50) = 5sqrt(2) diagonal of cube is from corner to corner this triangle has legs of 5 and 5sqrt(2)

OpenStudy (anonymous):

oh yeah... question said cube... sorry